Find the coordinates of the vertex for the parabola define by the given quadratic function:
f(x) = x to the second power + 8

Answer by oscargut(686)

in gral if f(x)=ax^2+bx+c then V=(-b/2a,f(-b/2a))
f(x)=x^2+8
V=(0,8)